calculate the voltage of a battery connected to a parallel plate capacitor with a plate ares of 3.0 cm2 and a plate separation of 5mm if the charge stored on the plates is 6.0 pC


1
Expert's answer
2021-04-26T17:04:58-0400

Find the capacitance:


"C=\frac{\epsilon_0A}{d}."

Find voltage:


"V=\frac QC=\frac{Qd}{\epsilon_0A}=\frac{(6·10^{-12})(0.005)}{(8.85·10^{-12})(3·10^{-6})}=1129\text{ V}."